Find the angle between normal to the planes (111) and (121) in a cubic unit cell. 


1
Expert's answer
2021-01-22T09:56:56-0500

"\alpha_1=90°-arctan\frac{\frac 11}{\frac 11}=90°-arctan1=90°-45°=45°."

"\alpha_2=90°-arctan\frac{\frac 12}{\frac11}=90°-arctan \frac 12=90°-27°=63°."

"\alpha=\alpha_2-\alpha_1=63°-45°=18°."